hhanan:

I know that 6+9=15 but i was not sure about the other numbers

hhanan:

Timmyspu:

There is a pattern going on here do you know what it is?

hhanan:

9 but for the other numbers are fractions

jhonyy9:

\[6*\frac{ 5 }{ 2 } = ?\]

jhonyy9:

\[15 *\frac{ 5 }{ 2 } = ?\]

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[6*\frac{ 5 }{ 2 } = ?\] \(\color{#0cbb34}{\text{End of Quote}}\) 15

jhonyy9:

and so much

jhonyy9:

yes 15 and the other one ?

hhanan:

37 1/2

Timmyspu:

I dont think its 15 because then you do that equation you end up getting 17.

jhonyy9:

6*5=30/2 = 15

Timmyspu:

ooooo ok my bad carry on.

jhonyy9:

6 time is there not 6 integer

hhanan:

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[15 *\frac{ 5 }{ 2 } = ?\] \(\color{#0cbb34}{\text{End of Quote}}\) and this is equal ?

jhonyy9:

15*5 = ?

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[15 *\frac{ 5 }{ 2 } = ?\] \(\color{#0cbb34}{\text{End of Quote}}\) and this is equal ? \(\color{#0cbb34}{\text{End of Quote}}\) 6

jhonyy9:

= 75 so \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[15 *\frac{ 5 }{ 2 } = ?\] \(\color{#0cbb34}{\text{End of Quote}}\) and this is equal ? \(\color{#0cbb34}{\text{End of Quote}}\) si this is equal \[\frac{ 75 }{ 2 }\]

jhonyy9:

from these result that to get the next term you need multiplie the last one by 5/2

hhanan:

thank you

jhonyy9:

anytime yw